OpenStudy (anonymous):

Solve each system by (a) the graphing method; (b) the substitution method. Pls Help 2)y=3x-7 6x=15y-1

OpenStudy (anonymous):

use y = 3x -7 and substitute into 6x = 15y -1 as such, 6x = 15(3x-7) - 1 then solve for x. using the value x you've got, substitute x-value into y = 3x -7 and get your y-value.